# Dynamical symmetry breaking in Georgi-Glashow chiral-gauge theories

### v1

@@ I Introduction
One of the outstanding challenges in Quantum Field Theory (QFT) and particle physics is to systematically understand the infrared (IR) phases of gauge theories in (3+1)(3+1)( 3 + 1 ) spacetime dimensions. A wide array of approaches, both exact and approximate, have been developed to probe the phases of such theories. Constraints have been developed from tools as the ’t Hooft anomaly matching conditions ’t Hooft (1980) or via connections to the supersymmetric limit Seiberg and Witten (1994a, b); Seiberg (1994, 1995). For vector-like gauge theories, the fermion path integral measure can be made real and positive by appropriately choosing fermion mass terms, thereby permitting the application of additional criteria such as the persistent mass condition [^1] See e.g. Ciambriello et al. (2024, 2025) for recent studies on the use of persistent mass conditions together with ’t Hooft anomaly matching conditions in a class of QCD-like theories, where it is proven that chiral symmetry must be spontaneously broken once the theory flows to an IR-free fixed point with full screening of color charges. Vafa and Witten (1984). Moreover, major insights into physical correlations in the strongly coupled regime have been obtained from first-principle computations such as Monte Carlo lattice simulations (see e.g. Wilson (1974); Kogut (1983); Greensite (2003); Bogolubsky et al. (2009); Bowman et al. (2004); Aoki et al. (2009)) and functional methods (see e.g. Alkofer and von Smekal (2001); Fischer (2006); Pawlowski et al. (2004); Braun et al. (2010); Cyrol et al. (2016); Aguilar et al. (2023)).
@@
Although most efforts have focused on gauge theories with vector-like fermions, due to its relevance to Quantum Chromodynamics (QCD) Gross et al. (2023), the study of chiral gauge theories remains of significant theoretical and phenomenological interest. On the one hand, a deeper understanding of chiral gauge theories is essential, as the Standard Model (SM) of particle physics is constructed based on a chiral matter content. However, strongly coupled chiral gauge theories Raby et al. (1980); Dimopoulos et al. (1980); Bars and Yankielowicz (1981) may potentially be relevant for addressing outstanding questions beyond the SM. However, our limited understanding of their strongly coupled behavior represents a major obstacle to assessing their phenomenological viability. Accordingly, it is of paramount importance to advance our knowledge of the dynamics governing these theories.
@@
By construction, no gauge-invariant fermion bilinear mass terms exist in chiral-gauge theories, which renders the persistent mass conditions inapplicable. Furthermore, the lattice regularization of chiral fermions remains a subtle and technically challenging task due to the Nielsen–Ninomiya theorem Nielsen and Ninomiya (1981a, b, c). These challenges have motivated a long-standing effort to develop robust theoretical tools capable of systematically probing the IR phases of chiral gauge theories. Recent progress has been made by exploiting ’t Hooft anomaly matching conditions associated with generalized global symmetries; for reviews, see e.g. Bolognesi et al. (2022, 2023); Konishi et al. (2024, 2024). Despite these advances, our understanding of chiral gauge dynamics remains incomplete, as either analytical or numerical methods are still largely unavailable.
@@
The IR phases of such theories are expected to depend crucially on the fermion content Sannino (2009); Poppitz and Unsal (2009); Bolognesi and Konishi (2019); Bolognesi et al. (2020, 2021, 2022, 2023, 2024). Some models may flow to an interacting IR fixed point, while others are conjectured to exhibit non-trivial Higgsing dynamics with condensate formation, leading to rich phenomena such as color-flavor locking, tumbling, and dynamical Abelianization. The standard lore is that one can smoothly interpolate between the Higgs and confinement regimes without encountering a phase transition Fradkin and Shenker (1979); Dimopoulos et al. (1980).
@@
In this paper, we study a class of anomaly-free and purely chiral gauge theories of the Georgi-Glashow (GG) type using the functional renormalization group (fRG) approach Wetterich (1993a); Dupuis et al. (2021). Due to its non-perturbative nature, the fRG is particularly suitable for investigating the strongly coupled dynamics of these theories. In this first work, we focus on the dynamical symmetry breaking (dSB) mechanisms which give rise to the highest condensate in the theory following in spirit previous studies in QCD-like theories Miransky (1999); Miransky and Yamawaki (1997); Gies and Jaeckel (2006); Braun and Gies (2006, 2010); Goertz et al. (2024); Braun et al. (2011). While here we focus on the start of dynamics, the present approach has proven its quantitative precision to describe dynamical chiral symmetry breaking and color confinement. Remarkably, it is shown in Cyrol et al. (2018); Mitter et al. (2015); Ihssen et al. (2024); Goertz et al. (2024); Fu et al. (2020) that the fRG approach can achieve quantitative precision with lattice results where they exist. Consequently, no conceptual complication is expected in the present approach.
@@
The remainder of the paper is organized as follows. In Section II we introduce the Georgi-Glashow model and its generalized form, which allows us to make contact to the loss of asymptotic freedom and a quantum scale-invariant regime. In Section III we introduce the effective action approach, the methodology for diagnosing dSB, and the fRG approach. In Section IV we discuss the results of the flow equation, concerning the resonant structure of the different channels and the phases in theory space, including the quantum scale invariant limit. Before concluding, we speculate about the IR fate of the theories and discuss the connection to other studies. Technical details concerning the basis construction and the flow equations as well as single examples can be found in the Appendices.
II The chiral theories
@@
where the fermion multiplicities are related in order to cancel the $[SU(N_{c})]^{3}$ triangle gauge anomaly. When $N_{c}=5$, the fermion content coincides with the model originally proposed by Georgi and Glashow in the context of grand unification of the SM gauge couplings Georgi and Glashow (1974). Due to this, models of this type with $N_{\textrm{gen}}=1$ are later referred to as chiral-gauge theories of the Georgi-Glashow type Raby et al. (1980); Dimopoulos et al. (1980), arguably the simplest purely chiral models.
@@
Theories with $N_{\textrm{gen}}<N_{\textrm{gen}}^{\rm AF}=(11N_{c})/(2N_{c}-6)$ have a well-defined asymptotically free ultraviolet (UV) limit. When $N_{\textrm{gen}}$ is slightly below $N_{\textrm{gen}}^{\rm AF}$, it is believed that theories flow into an interacting IR fixed point, analogous in spirit to the Caswell-Banks-Zaks Caswell (1974); Banks and Zaks (1982) solutions in QCD-like theories. However, unlike the latter, determining the existence of the fixed point in the GG-type models is tricky. In particular, a perturbative analysis of the existence of a fixed point is not valid, since the zeros of the $\beta$ function cannot be parametrically under control [^2] It may still be numerically small, but this is not sufficient to prove the existence of an interacting fixed point in the perturbative regime., contrary to QCD-like theories in the Veneziano-Witten limit. A simple estimation using the leading two coefficients of the $\beta$ function suggests that all the interacting fixed points, if they exist at all, are non-perturbative, even when the gauge coupling can be numerically small; see Bolognesi et al. (2023) for a review emphasizing this point. For sufficiently small $N_{\textrm{gen}}$, fixed-point solutions are not found and the gauge coupling drives the appearance of so-far unknown dynamics which we target in this work.
We note that while the strongly coupled regime of the one-generation model was studied in Raby et al. (1980); Dimopoulos et al. (1980) long ago, the multi-generation GG models, which we refer to as the generalized, have received very little attention; for instance, see only Bolognesi and Konishi (2019); Bai and Stolarski (2022) for some recent studies.
III Effective action and the functional approach
@@
The dynamical emergence of fermion condensates has a long history in the context of effective actions, particularly in condensed matter and nuclear physics Nambu and Jona-Lasinio (1961); Braun (2012). This mechanism is encoded in the fermionic $n$-point functions and can commonly be diagnosed in a divergence of the dressings. Through a Hubbard–Stratonovich transformation Hubbard (1959); Stratonovich (1957), this divergence is shown to be mathematically equivalent to the appearance of a condensate in the bosonized formulation of the theory. In that setting, the formation of a condensate corresponds to a change in the curvature of the bosonic effective potential around the origin, specifically, a change in the sign of the boson mass term. Thus, the transition from a symmetric to a broken phase is indicated by a vanishing boson mass (i.e., an infinite correlation length), or equivalently, by a divergence in the corresponding four-fermion coupling. See, e.g., Miransky and Yamawaki (1989); Miransky (1994); Gies and Jaeckel (2006) for illustrative examples.
In this work, we follow these lines to analyze the appearance of condensates in generalized GG models. To account for quantum corrections and derive the $n$-point functions of the effective action, we employ the functional renormalization group Wetterich (1993a). In this non-perturbative approach, momentum modes in the path integral measure are integrated out progressively with the help of an IR regulator function $R_{k}$, which suppresses fluctuations below the cutoff scale $k$. This procedure defines the effective average action $\Gamma_{k}$ Wetterich (1991, 1993b), which incorporates quantum fluctuations down to the momentum scale $k$ and recovers the full effective action in the limit $k\to 0$. Central to this approach is the flow or Wetterich equation with which we implement quantum corrections in a non-perturbative manner,
@@
Exactly solving the flow equation (3) is generally not feasible, and practical applications rely on truncations of the effective average action. These consist of considering a finite set of operators consistent with the symmetries of the theory, each with scale-dependent dressings. In this work, we consider the minimal truncation including up to four-fermion operators which permit diagnosing spontaneous symmetry breaking. The regulator function $R_{k}$ is not unique and can be chosen to optimize convergence and numerical stability. We employ the Litim regulator Litim (2001) which provides analytic handle on the flow equations. We refer technical details on the flow equation and truncation to Appendix B.
Functional approaches are widely used to study dSB, particularly in QCD and its phase structure, which are close in spirit to the theories studied here. Advanced fRG computations incorporate explicit momentum dependence in higher $n$-point functions, both in vacuum and at finite temperature or chemical potential, see e.g. Fu et al. (2023, 2024, 2025); Braun et al. (2018, 2020); Ihssen et al. (2024). Moreover, the fRG framework allows for scale-dependent field transformations along the flow, enabling the description of dynamically emerging effective degrees of freedom Gies and Wetterich (2004); Fukushima et al. (2021); Pawlowski (2007). The emergent composites (or dynamical hadronisation) approach facilitates the inclusion of higher-dimensional fermionic operators via bosonic interactions $\phi^{n}\propto\left(\bar{\psi}\,{\cal T}\psi\right)^{n}$, thus extending the analysis beyond the singularity of the four-fermion coupling that signals dSB and providing smooth access to the chirally broken regime. This formulation offers both technical and conceptual advantages for studying strongly coupled gauge–fermion systems.
In this first work, we restrict ourselves to a qualitative analysis and leave the full bosonized treatment to future studies. Our goal is to identify which condensate forms at the highest scales, which is the first step towards the full IR dynamics. For this, it suffices to employ a minimal Fierz-complete truncation of four-fermion operators Gies and Jaeckel (2006); Braun and Gies (2007, 2006); Braun et al. (2011).
@@
$\Gamma_{k}\left[A_{\mu},\bar{c},c,\bar{\psi},\psi,\bar{\chi},\chi\right]=$ $\,\Gamma_{{\rm gauge},k}\left[A_{\mu},\bar{c},c\right]$ (4)
$\hskip-99.58464pt+\Gamma_{{\rm Dirac},k}\left[A_{\mu},\bar{\psi},\psi,\bar{\chi},\chi\right]+\Gamma_{{\rm 4F},k}\left[\bar{\psi},\psi,\bar{\chi},\chi\right]+\ldots\,,$
@@
$\Gamma_{{\rm 4F},k}[\bar{\psi},\psi,\bar{\chi},\chi]=$ (5)
$\quad-\int_{x}Z_{\psi}^{2}\,\sum_{i=1}^{2}\lambda_{i}{\cal O}_{i}+Z_{\chi}^{2}\sum_{i=3}^{5}\lambda_{i}{\cal O}_{i}+Z_{\psi}Z_{\chi}\sum_{i=6}^{7}\lambda_{i}{\cal O}_{i}\,,$
and include all possible four-fermion operators forming a Fierz-complete basis which we construct with the Young Tensor Method Li et al. (2021a, b, 2022). In (5), $Z_{\psi}$ and $Z_{\chi}$ are the fermionic field wave functions and $\lambda_{i}=k^{-2}\bar{\lambda}_{i}$ are the dimensionless form of the four-fermion dressings for the set of operators
@@
IV Patterns of symmetry breaking
The evolution of four-fermion dressings is dictated by their respective RG flows, which are triggered purely by the gauge dynamics. While these flows have been extensively studied in QCD-like theories (see e.g. Goertz et al. (2024); Fu et al. (2023, 2024, 2025); Braun (2012)), as we will see, the situation in chiral-gauge theories with several species is qualitatively different and remains rather unexplored. We commence this Section by examining the general structure of the four-fermion RG flow in gauge–fermion systems and the underlying driving mechanism. We then discuss the resonant structure in this class of theories and provide more details about particular theories in Appendix C.

### v2

@@ 1 Introduction
One of the outstanding challenges in Quantum Field Theory (QFT) and particle physics is to systematically understand the infrared (IR) phases of gauge theories in $(3+1)$ spacetime dimensions. A wide array of approaches, both exact and approximate, have been developed to probe the phases of such theories. Constraints have been developed from tools as the ’t Hooft anomaly matching conditions tHooft:1979rat or via connections to the supersymmetric limit Seiberg:1994rs ; Seiberg:1994aj ; Seiberg:1994bz ; Seiberg:1994pq . For vector-like gauge theories, the fermion path integral measure can be made real and positive by appropriately choosing fermion mass terms, thereby permitting the application of additional criteria such as the persistent mass condition [^1] See e.g. Ciambriello:2022wmh ; Ciambriello:2024xzd for recent studies on the use of persistent mass conditions together with ’t Hooft anomaly matching conditions in a class of QCD-like theories, where it is proven that chiral symmetry must be spontaneously broken once the theory flows to an IR-free fixed point with full screening of color charges. Vafa:1983tf . Moreover, major insights into physical correlations in the strongly coupled regime have been obtained from first-principle computations such as Monte Carlo lattice simulations (see e.g. Wilson:1974sk ; Kogut:1982ds ; Greensite:2003bk ; Bogolubsky:2009dc ; Bowman:2004jm ; Aoki:2009sc ) and functional methods (see e.g. Alkofer:2000wg ; Fischer:2006ub ; Pawlowski:2003hq ; Braun:2007bx ; Cyrol:2016tym ; Aguilar:2022thg ).
@@
Although most efforts have focused on gauge theories with vector-like fermions, due to its relevance to Quantum Chromodynamics (QCD) Gross:2022hyw , the study of chiral gauge theories remains of significant theoretical and phenomenological interest. On the one hand, a deeper understanding of chiral gauge theories is essential, as the Standard Model (SM) of particle physics is constructed based on a chiral matter content. However, strongly coupled chiral gauge theories Raby:1979my ; Dimopoulos:1980hn ; Bars:1981se may potentially be relevant for addressing outstanding questions beyond the SM. However, our limited understanding of their strongly coupled behavior represents a major obstacle to assessing their phenomenological viability. Accordingly, it is of paramount importance to advance our knowledge of the dynamics governing these theories.
@@
By construction, no gauge-invariant fermion bilinear mass terms exist in chiral-gauge theories, which renders the persistent mass conditions inapplicable. Furthermore, the lattice regularization of chiral fermions remains a subtle and technically challenging task due to the Nielsen–Ninomiya theorem Nielsen:1980rz ; Nielsen:1981hk ; Nielsen:1981xu . These challenges have motivated a long-standing effort to develop robust theoretical tools capable of systematically probing the IR phases of chiral gauge theories. Recent progress has been made by exploiting ’t Hooft anomaly matching conditions associated with generalized global symmetries; for reviews, see e.g. Bolognesi:2021jzs ; Bolognesi:2023sxe ; Konishi:2024rjz . Despite these advances, our understanding of chiral gauge dynamics remains incomplete, as either analytical or numerical methods are still largely unavailable.
@@
The IR phases of such theories are expected to depend crucially on the fermion content Sannino:2009za ; Poppitz:2009uq ; Bolognesi:2019wfq ; Bolognesi:2020mpe ; Bolognesi:2021yni ; Bolognesi:2021jzs ; Bolognesi:2023sxe ; Bolognesi:2024bnm . Some models may flow to an interacting IR fixed point, while others are conjectured to exhibit non-trivial Higgsing dynamics with condensate formation, leading to rich phenomena such as color-flavor locking, tumbling, and dynamical Abelianization. The standard lore is that one can smoothly interpolate between the Higgs and confinement regimes without encountering a phase transition Fradkin:1978dv ; Dimopoulos:1980hn .
@@
In this paper, we study a class of anomaly-free and purely chiral gauge theories of the Georgi-Glashow (GG) type using the functional renormalization group (fRG) approach Wetterich:1992yh ; Dupuis:2020fhh . Due to its non-perturbative nature, the fRG is particularly suitable for investigating the strongly coupled dynamics of these theories. In this first work, we focus on the dynamical symmetry breaking (dSB) mechanisms which give rise to the highest condensate in the theory following in spirit previous studies in QCD-like theories Miransky:1998dh ; Miransky1997 ; Gies:2005as ; Braun:2006jd ; Braun:2009ns ; Goertz:2024dnz ; Braun:2010qs . While here we focus on the start of dynamics, the present approach has proven its quantitative precision to describe dynamical chiral symmetry breaking and color confinement. Remarkably, it is shown in Cyrol:2017ewj ; Mitter:2014wpa ; Ihssen:2024miv ; Goertz:2024dnz ; Fu:2019hdw that the fRG approach can achieve quantitative precision with lattice results where they exist. Consequently, no conceptual complication is expected in the present approach.
@@
The remainder of the paper is organized as follows. In Section 2 we introduce the Georgi-Glashow model and its generalized form, which allows us to make contact to the loss of asymptotic freedom and a quantum scale-invariant regime. In Section 3 we introduce the effective action approach, the methodology for diagnosing dSB, and the fRG approach. In Section 4 we discuss the results of the flow equation, concerning the resonant structure of the different channels and the phases in theory space, including the quantum scale invariant limit. Before concluding, we speculate about the IR fate of the theories and discuss the connection to other studies. Technical details concerning the basis construction and the flow equations as well as single examples can be found in the Appendices.
2 The chiral theories
@@
where the fermion multiplicities are related in order to cancel the $SU(N_{c})]^{3}$ triangle gauge anomaly. When $N_{c}=5$, the fermion content coincides with the model originally proposed by Georgi and Glashow in the context of grand unification of the SM gauge couplings Georgi:1974sy . Due to this, models of this type with $N_{\textrm{gen}}=1$ are later referred to as chiral-gauge theories of the Georgi-Glashow type Raby:1979my ; Dimopoulos:1980hn , arguably the simplest purely chiral models.
@@ 11 $(N_{c}-2)$
Theories with $N_{\textrm{gen}}<N_{\textrm{gen}}^{\rm AF}=(11N_{c})/(2N_{c}-6)$ have a well-defined asymptotically free ultraviolet (UV) limit. When $N_{\textrm{gen}}$ is slightly below $N_{\textrm{gen}}^{\rm AF}$, it is believed that theories flow into an interacting IR fixed point, analogous in spirit to the Caswell-Banks-Zaks Caswell:1974gg ; Banks:1981nn solutions in QCD-like theories. However, unlike the latter, determining the existence of the fixed point in the GG-type models is tricky. In particular, a perturbative analysis of the existence of a fixed point is not valid, since the zeros of the $\beta$ function cannot be parametrically under control [^2] It may still be numerically small, but this is not sufficient to prove the existence of an interacting fixed point in the perturbative regime., contrary to QCD-like theories in the Veneziano-Witten limit. A simple estimation using the leading two coefficients of the $\beta$ function suggests that all the interacting fixed points, if they exist at all, are non-perturbative, even when the gauge coupling can be numerically small; see Bolognesi:2023sxe for a review emphasizing this point. For sufficiently small $N_{\textrm{gen}}$, fixed-point solutions are not found and the gauge coupling drives the appearance of so-far unknown dynamics which we target in this work.
We note that while the strongly coupled regime of the one-generation model was studied in Raby:1979my ; Dimopoulos:1980hn long ago, the multi-generation GG models, which we refer to as the generalized, have received very little attention; for instance, see only Bolognesi:2019wfq ; Bai:2021tgl for some recent studies.
3 Effective action and the functional approach
@@
The dynamical emergence of fermion condensates has a long history in the context of effective actions, particularly in condensed matter and nuclear physics Nambu:1961fr ; Braun:2011pp . This mechanism is encoded in the fermionic $n$-point functions and can commonly be diagnosed in a divergence of the dressings. Through a Hubbard–Stratonovich transformation HubbardPhysRevLett.3.77 ; Stratonovich , this divergence is shown to be mathematically equivalent to the appearance of a condensate in the bosonized formulation of the theory. In that setting, the formation of a condensate corresponds to a change in the curvature of the bosonic effective potential around the origin, specifically, a change in the sign of the boson mass term. Thus, the transition from a symmetric to a broken phase is indicated by a vanishing boson mass (i.e., an infinite correlation length), or equivalently, by a divergence in the corresponding four-fermion coupling. See, e.g., Miransky1989b ; Miransky:1994vk ; Gies:2005as for illustrative examples.
In this work, we follow these lines to analyze the appearance of condensates in generalized GG models. To account for quantum corrections and derive the $n$-point functions of the effective action, we employ the functional renormalization group Wetterich:1992yh . In this non-perturbative approach, momentum modes in the path integral measure are integrated out progressively with the help of an IR regulator function $R_{k}$, which suppresses fluctuations below the cutoff scale $k$. This procedure defines the effective average action $\Gamma_{k}$ Wetterich:1989xg ; Wetterich:1991be , which incorporates quantum fluctuations down to the momentum scale $k$ and recovers the full effective action in the limit $k\to 0$. Central to this approach is the flow or Wetterich equation with which we implement quantum corrections in a non-perturbative manner,
@@
Exactly solving the flow equation (3) is generally not feasible, and practical applications rely on truncations of the effective average action. These consist of considering a finite set of operators consistent with the symmetries of the theory, each with scale-dependent dressings. In this work, we consider the minimal truncation including up to four-fermion operators which permit diagnosing spontaneous symmetry breaking. The regulator function $R_{k}$ is not unique and can be chosen to optimize convergence and numerical stability. We employ the Litim regulator Litim:2001up which provides analytic handle on the flow equations. We refer technical details on the flow equation and truncation to Appendix B.
Functional approaches are widely used to study dSB, particularly in QCD and its phase structure, which are close in spirit to the theories studied here. Advanced fRG computations incorporate explicit momentum dependence in higher $n$-point functions, both in vacuum and at finite temperature or chemical potential, see e.g. Fu:2022uow ; Fu:2024ysj ; Fu:2025hcm ; Braun:2018bik ; Braun:2019aow ; Ihssen:2024miv . Moreover, the fRG framework allows for scale-dependent field transformations along the flow, enabling the description of dynamically emerging effective degrees of freedom Gies:2002hq ; Fukushima:2021ctq ; Pawlowski:2005xe . The emergent composites (or dynamical hadronisation) approach facilitates the inclusion of higher-dimensional fermionic operators via bosonic interactions $\phi^{n}\propto\left(\bar{\psi}\,{\cal T}\psi\right)^{n}$, thus extending the analysis beyond the singularity of the four-fermion coupling that signals dSB and providing smooth access to the chirally broken regime. This formulation offers both technical and conceptual advantages for studying strongly coupled gauge–fermion systems.
In this first work, we restrict ourselves to a qualitative analysis and leave the full bosonized treatment to future studies. Our goal is to identify which condensate forms at the highest scales, which is the first step towards the full IR dynamics. For this, it suffices to employ a minimal Fierz-complete truncation of four-fermion operators Gies:2005as ; Braun:2005uj ; Braun:2006jd ; Braun:2010qs .
@@
$\Gamma_{k}\left[A_{\mu},\bar{c},c,\bar{\psi},\psi,\bar{\chi},\chi\right]=$ $\,\Gamma_{{\rm gauge},k}\left[A_{\mu},\bar{c},c\right]+\Gamma_{{\rm Dirac},k}\left[A_{\mu},\bar{\psi},\psi,\bar{\chi},\chi\right]+\Gamma_{{\rm 4F},k}\left[\bar{\psi},\psi,\bar{\chi},\chi\right]+\ldots\,$ (4)
@@
$\Gamma_{{\rm 4F},k}[\bar{\psi},\psi,\bar{\chi},\chi]=$ $-\int_{x}Z_{\psi}^{2}\,\sum_{i=1}^{2}\lambda_{i}{\cal O}_{i}+Z_{\chi}^{2}\sum_{i=3}^{5}\lambda_{i}{\cal O}_{i}+Z_{\psi}Z_{\chi}\sum_{i=6}^{7}\lambda_{i}{\cal O}_{i}\,$, (5)
and include all possible four-fermion operators forming a Fierz-complete basis which we construct with the Young Tensor Method Li:2020gnx ; Li:2020xlh ; Li:2022tec . In (5), $Z_{\psi}$ and $Z_{\chi}$ are the fermionic field wave functions and $\bar{\lambda}_{i}=k^{2}\;\lambda_{i}$ are the dimensionless form of the four-fermion dressings for the set of operators,
@@
4 Patterns of symmetry breaking
The evolution of four-fermion dressings is dictated by their respective RG flows, which are triggered purely by the gauge dynamics. While these flows have been extensively studied in QCD-like theories (see e.g. Goertz:2024dnz ; Dupuis:2020fhh ; Fu:2022uow ; Fu:2024ysj ; Fu:2025hcm ; Braun:2011pp ), as we will see, the situation in chiral-gauge theories with several species is qualitatively different and remains rather unexplored. We commence this Section by examining the general structure of the four-fermion RG flow in gauge–fermion systems and the underlying driving mechanism. We then discuss the resonant structure in this class of theories and provide more details about particular theories in Appendix C.
Figure 1: Sketch of the diagram topologies entering the RG flow of the four-$\psi$ point function relevant to the flow of the four-fermion couplings in (8). Plain arrowed lines represent full $\psi$ propagators, dashed lines the respective $\chi$ propagators, and curly lines the full gauge boson propagators. The operator $\tilde{\partial}_{t}$ denotes a shorthand for the insertion of the regulators ($\partial_{t}R_{k}$) inherent to the flow equation (3), avoiding the need to write the different combinations of regulated lines explicitly. The $\ldots$ indicate other diagrams involving other higher dimensional operators (e.g. six-point functions), which are neglected in this analysis.